### v2

# Rethinking Optimal Verification Granularity for Compute-Efficient Test-Time Scaling

Abstract
Test-time scaling (TTS) has proven effective in enhancing the reasoning capabilities of large language models (LLMs). Verification plays a key role in TTS, simultaneously influencing (1) reasoning performance and (2) compute efficiency, due to the quality and computational cost of verification. In this work, we challenge the conventional paradigms of verification, and make the first attempt toward systematically investigating the impact of verification granularity—that is, how frequently the verifier is invoked during generation, beyond verifying only the final output or individual generation steps. To this end, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes beam search and Best-of-N sampling via a tunable granularity parameter $g$. Extensive experiments with VG-Search under varying compute budgets, generator-verifier configurations, and task attributes reveal that dynamically selecting $g$ can improve the compute efficiency and scaling behavior. Building on these findings, we propose adaptive VG-Search strategies that achieve accuracy gains of up to 3.1% over Beam Search and 3.6% over Best-of-N, while reducing FLOPs by over 52%. Our code is avaiblae at github.com/hmarkc/VG-Search.
1 Introduction
The past few years have witnessed the rapid advancement of large language models (LLMs), driven by scaling of model size and training data [6, 17, 28]. However, further training-time scaling is increasingly constrained by prohibitive computational costs and the limited availability of high-quality human-generated data [34]. Test-time scaling (TTS) [33, 5, 32] offers a promising alternative by enhancing performance through additional compute at inference time. TTS techniques generally fall into two categories: internal scaling [33, 25], which focuses on optimizing a single generation trajectory, and sampling-based scaling [5, 32, 35], which improves performance by exploring multiple candidate generations. These two approaches are orthogonal and can be combined to achieve higher performance [22]. Sampling-based methods are typically training-free and easy to integrate with internal scaling approaches in a plug-and-play manner.
Verification, commonly implemented as a learned reward model or scoring function, plays a key role in both TTS paradigms. State-of-the-art (SOTA) sampling-based methods, such as Diverse Verifier Tree Search (DVTS) [4] and verifier-guided Beam Search [32], leverage a separate verifier LLM to guide the generation of a generator LLM, improving sampling efficiency and accuracy. In these methods, a generation step is typically defined as a chunk of text delimited by special tokens (e.g., newline characters), which serves as the atomic unit of verification. However, this verification granularity choice is heuristic [32, 22] and remains static [36], with no guarantee of optimality. As shown in Figure 1(a), our profiling results show that verifier scores often remain stable across multiple generation steps (e.g., over 50% of 2-step score differences are below $1\%$ of score range), suggesting redundancy in the current verification granularity. This inefficiency causes verification to account for an increasingly large proportion of the overall inference latency, as illustrated in Figure 1(b). These insights and observations motivate us to explore the following two research questions (RQs):
(a) Score Stability
(b) Verification Cost
(c) Inference Latency
(d) Optimized Performance
Figure 1: Motivations for Optimizing Verification Granularity $g$. (a) Small PRM score deltas (LLaMA3.1-8B-PRM) across generation steps indicate redundant verification. (b) Verification incurs notable latency overheads. (c) Increasing verification granularity lowers the latency of both Proposer and Verifier. (d) Optimizing $g$ improves the accuracy-compute tradeoff over fixed-granularity baselines (Qwen2.5-Math-7B as generator, Skywork-o1-1.5B as verifier).
  • RQ1. Is the conventional verification granularity optimal for the accuracy–compute scaling?
  • RQ2. If not, how can we optimize it to achieve a better accuracy–compute frontier?
To explore RQ1, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes verifier-guided Beam Search and Best-of-N methods. VG-Search employs a granularity parameter $g$ that controls the verification granularity (Section 2). Our hardware performance profiling (Figure 1(c)) shows that $g$ has a significant impact on the latency contributions of both the generator LLM and the verifier. To further investigate the accuracy–compute trade-off under varying $g$, we conduct extensive experiments with VG-Search across different values of $g$, models, compute budget, and datasets (Section 4). Building on top of the insights from RQ1, we propose an adaptive VG-Search approach which dynamically adjusts the verification granularity based on the compute budget and task attributes (Section 5) to further address RQ2. As shown in Figure 1(d), our results demonstrate better performance–compute frontiers than previous sampling-based TTS methods. Overall, our main contributions are:
  • We propose Variable Granularity Search (VG-Search), enabling systematic analysis of how verification granularity affects performance and compute scaling.
  • Through extensive experiments, we show that the conventional static verification granularity is sub-optimal, limiting the performance scaling of verifier-guided search methods.
  • We develop adaptive VG-Search strategies that dynamically tune granularity, achieving up to 3.1% higher accuracy than Beam Search, 3.6% over Best-of-N, with over 52% compute reduction.
2 Exploring Verification Granularity for Reasoning
2.1 Verification Granularity in Verifier-Guided Search
In verifier-guided search, a generator model proposes candidate steps and a verifier model evaluates them to guide the search process. The task of generating a correct solution can be framed as a collaboration between two components:
  • Generator (Proposer) $\mathcal{G}$: A probabilistic model, typically an LLM, responsible for generating a partial or complete candidate sequence $\mathbf{s}_{t:t+g}$, where $g$ denotes the number of the generation steps. The generation process is governed by the distribution $P(\mathbf{s}_{t:t+g}\mid\mathcal{G})$.
  • Verifier $\mathcal{V}$: A scoring function $\mathcal{V}(\mathbf{s}_{t:t+g})\rightarrow\mathbb{R}$ that assigns a scalar score to each partial or complete candidate sequence $\mathbf{s}_{t:t+g}$. The verifier is invoked after every $g$ generation steps. This score is intended to correlate with the expected quality or correctness of the sequence prefix.
The goal of the search process is to generate a correct or high-quality solution $\mathbf{s}_{1:T}^{*}$ of length $T$ via interaction between the generator and verifier. A key hyperparameter in this process is the verification granularity $g$, which controls how frequently the verifier can intervene and guide the generator. To illustrate the impact of $g$, we link the verifier-guided search to an edge case, the Infinite Monkey Theorem (IMT) [38], where a random generator (monkey) tries to reproduce a target sequence (e.g., Shakespeare’s work) under a perfect verifier invoked every $g$ characters. When verification is infrequent (large $g$), the expected number of attempts grows exponentially as $\mathbb{E}[A^{g}]$, making success computationally infeasible. In contrast, with frequent verification, early error pruning reduces the expected cost $\mathbb{E}[A^{g}]$, at the expense of higher verification overhead. This toy setting underscores a fundamental trade-off between verification cost and exploration efficiency. Optimizing $g$ is thus critical for effective collaboration between generator $\mathcal{G}$ and verifier $\mathcal{V}$ under compute constraints.
The most common approach to defining $g$ relies on heuristics and remains static. However, this convention might be neither consistent nor optimal: the model distribution varies widely—some generators are more verbose, others more structured—and the number of tokens required to conceptualize a coherent "thinking step" can differ significantly across tasks. This motivates a systematic study of how performance scales with different granularity $g$.
Figure 2: Variable-Granularity Search (VG-Search). VG-Search unifies verifier-guided Beam Search and Best-of-N via the granularity parameter $g$. At each step, $B_{1}\times B_{2}$ candidates are verified and $B_{1}=2$ candidates are selected. $B_{2}=2$ continuations are generated per selected path. Only the $B_{1}$ selected candidates proceed to the Extend Step, reducing proposer FLOPs by pruning early.
2.2 Variable-Granularity Search
Beam Search [5] and Best-of-N sampling [32] are two prevalent methods for scaling large language model (LLM) performance at test-time by leveraging additional computation. These methods represent two opposing extremes within the spectrum of verifier-guided search, primarily differing in the frequency with which the verifier $\mathcal{V}$ checks the outputs from the proposer $\mathcal{G}$. In Beam Search, verification is highly fine-grained: $\mathcal{V}$ actively guides candidate selection at each generation step, enabling the early pruning of less promising paths. Conversely, Best-of-N sampling employs a coarse-grained approach, where the verification is performed on an entire candidate solution. While this allows for broader exploration, it lacks the intermediate guidance provided by frequent verification.
To bridge this gap and enable a continuous exploration of verification granularity, we introduce Variable-Granularity Search (VG-Search). VG-Search unifies Beam Search and Best-of-N sampling through the concept of verification granularity ($g$). By varying $g$, VG-Search allows for continuous control over the trade-off between fine-grained guidance and broad exploration. Key hyperparameters of VG-Search include:
  • Beam Width ($B_{1}$): The number of top-scoring candidate sequences (beams) retained after verification and selection.
  • Branch-Out Factor ($B_{2}$): The number of alternative continuations generated by $\mathcal{G}$ from each of the $B_{1}$ retained beams before the next verification phase.
  • Verification Granularity ($g$): The number of generation steps evaluated per verification—i.e., the interval between verifier ($\mathcal{V}$) calls.
As illustrated in Figure 2, a single cycle of VG-Search advances the search from a candidate length $t$ to $t+g$. The full algorithm proceeds as follows:
  1. 1.
    Start: Initialize with $B_{1}\times B_{2}$ candidates from the initial prompt.
  2. 2.
    Verify & Select: Evaluate $B_{1}\times B_{2}$ candidates using $\mathcal{V}$, and retain the top $B_{1}$ beams.
  3. 3.
    Extend: For each of the $B_{1}$ selected beams, produce g1g-1 generation steps using $\mathcal{G}$.
  4. 4.
    Branch: For each extended beam, produce $B_{2}$ single-generation-step continuations using $\mathcal{G}$.
  5. 5.
    Repeat: Go back to Step 2 and iterate until termination criteria are met.
Omitting Step 3 (Extend) or setting $g=1$ reduces VG-Search to standard Beam Search. At the other extreme, when $g=L$ (the number of generation steps in a full solution), VG-Search becomes equivalent to Best-of-N sampling. One alternative design is to produce g1g-1 generation steps in Branch, mimicking Beam Search but with variable verification granularity. However, prior work [32] shows that rolling out multiple future tokens, as in LookAhead Search, does not yield the best performance. We therefore adopt a simpler, more compute-efficient design: Branch generates just one step per candidate, and early pruning via the verifier occurs before Extend. As a result, only the top $B_{1}$ candidates continue in Extend, instead of the full $B_{1}\times B_{2}$ set used in Beam Search or Best-of-N—leading to significantly lower proposer FLOPs. Larger $g$ further amplifies savings for both verifier and proposer. We validate this design in Section 4.1. A case study of VG-Search is illustrated in Appendix A.8.
2.3 Compute Cost Model for Generator and Verifier
To analyze and optimize $g$, we define a compute cost model in FLOPs with the following parameters: (1) $L$: average solution length in generation steps, (2) $P_{g}$, $P_{v}$: parameter counts of the generator and verifier, (3) $F_{\text{g}}$: FLOPs per generation step, (4) $F_{\text{v}}$: FLOPs per verifier call. The number of VG-Search cycles is approximately $N_{\text{cycles}}=L/g$. Following prior work [31], we approximate $F_{\text{g}}\approx 2TP_{g}$, where $T$ is the average number of tokens per generation step. Similarly, $F_{\text{v}}\approx 2\alpha P_{v}$, where $\alpha$ is a verifier-specific variable.
The total compute for generation, $C_{\mathcal{G}}$, includes extending the $B_{1}$ primary beams and generating $B_{2}$ branches across all cycles:
$C_{\mathcal{G}}=B_{1}\cdot\left(\underbrace{g-1}_{\text{Extending}}+\underbrace{B_{2}}_{\text{Branching}}\right)\cdot N_{\text{cycles}}\cdot F_{\text{g}}=2\cdot T\cdot B_{1}\cdot(g-1+B_{2})\cdot\frac{L}{g}\cdot P_{g}$ (1)
The total compute for verification, $C_{\mathcal{V}}$, accounts for scoring $B_{1}\cdot B_{2}$ candidates across all cycles:
$C_{\mathcal{V}}=B_{1}\cdot B_{2}\cdot N_{\text{cycles}}\cdot F_{\text{v}}=2\cdot\alpha\cdot B_{1}\cdot B_{2}\cdot\frac{L}{g}\cdot P_{v}$ (2)
The total compute cost $C_{\text{total}}$ becomes:
$C_{\text{total}}=C_{\mathcal{G}}+C_{\mathcal{V}}=2\cdot B_{1}\cdot\frac{L}{g}\cdot\left[T\cdot(g-1+B_{2})\cdot P_{g}+\alpha\cdot B_{2}\cdot P_{v}\right]$ (3)
We adopt the proposed cost model to study the scaling behavior of VG-Search (Section 4), which provides a reliable proxy for measured inference latency as demonstrated in Figure 1(c).
3 Experiment Setup
To investigate the optimality of the conventional verification granularity mentioned in RQ1, we conduct experiments on mathematical reasoning benchmarks.
Task and Datasets. We evaluate on the MATH-500 [21] and AIME [2] datasets. MATH-500 samples 500 problems from the MATH [16] benchmark, with difficulty levels labelled. AIME comprises 90 advanced high school mathematics problems from the past three years (AIME24, AIME23, AIME22). Additionally, we sample 250 problems from the MATH [16] benchmark as a validation set for Section 5, referred to as MATH-250. It consists of 50 problems per difficulty level.
[…]
4 Understanding the Limits of the Verification Granularity Convention
4.1 Test-Time Scaling Law with Verification Granularity
In this section, we aim to answer RQ1, investigating how the optimal verification granularity $g$ varies with both the compute budget and the capabilities of the Proposer-Verifier pair. To explore different operating points on the accuracy-compute trade-off curve (Figure 3), we vary $g$ and adjust the number of generations $n$ to modulate the total compute budget. We analyze three representative generator ($\mathcal{G}$)–verifier ($\mathcal{V}$) pairs, specifically chosen to probe the interplay between model capability and compute allocation:
  1. 1.
    Strong $\mathcal{G}$ (Qwen2.5-Math-7B) with a Small $\mathcal{V}$ (Skywork-o1-1.5B),
  2. 2.
    Weak $\mathcal{G}$ (Llama-3.2-3B-Instruct, not math-finetuned) with a Small $\mathcal{V}$ (Skywork-o1-1.5B),
  3. 3.
    Medium $\mathcal{G}$ (Qwen2.5-Math-1.5B) with a Large $\mathcal{V}$ (Skywork-o1-7B).
Figure 3: Accuracy vs. compute (FLOPs) across different verification granularities $g$.
Our findings in Figure 3 indicate that the optimal verification granularity $g$ depends on both the generator’s capability and the available compute budget:
Strong Generators Prefer Sparse Verification, while weak ones need frequent checks. For the strong Qwen2.5-Math-7B generator paired with a small verifier, while standard Beam Search ($g=1$) is effective at lower compute costs, sparser verification ($g\in\{2,3,4\}$) achieves superior accuracy at medium to high compute budgets, as shown in the first column of Figure 3. Notably, $g=3$ reaches the highest peak accuracy, outperforming $g=1$ by approximately 4% on MATH-500. This suggests that a strong generator can reliably produce longer correct partial solutions, making frequent verification less critical and allowing compute to be reallocated (e.g., to wider beams via $B_{1}$) for better overall performance. As shown in the second column of Figure 3, moderately strong Qwen2.5-Math-1.5B, when paired with a large verifier, also benefits from sparser verification, with $g=2$ consistently outperforming $g=1$ from mid-to-high compute and achieving the best peak accuracy. However, excessively sparse verification ($g=4$) becomes detrimental, indicating an optimal balance point. In contrast, as shown in the last column of Figure 3, the non-specialized Llama-3B generator has peak performance with frequent verification ($g=1$) on MATH-500 and MATH-250, implying that weaker or less task-aligned generators require continuous guidance. Interestingly, on the more challenging AIME dataset, all Proposer-Verifier pairs show a tendency for sparser verification ($g>1$, sometimes even $g=4$) to outperform $g=1$ at medium and high compute budgets. This might be because AIME problems benefit more from the broader exploration encouraged by sparser verification (which makes the search resemble Best-of-N), or the PRM’s step-wise utility is less consistently high on AIME, reducing the benefit of frequent checks.
Optimal Granularity Varies with Compute Budget. Across most models and datasets, sparser verification ($g>1$) tends to become more competitive as the total compute budget increases. At very low compute budgets, the aggressive pruning of standard Beam Search ($g=1$) often provides a more robust performance baseline. This aligns with the intuition that investing in more extended generation phases between verifications is more viable when more overall compute is available to support wider exploration or more candidate paths.
Optimal Granularity Saves Computation Significantly. A key advantage of using sparser verification ($g>1$) is the potential for substantial FLOPs savings while maintaining—or even improving—performance. For example, in the Strong $\mathcal{G}$, Small $\mathcal{V}$ setting on MATH-500, setting $g=3$ achieves approximately 88% accuracy at $\sim 2^{13}$ FLOPs, whereas $g=1$ requires $\sim 2^{15}$ FLOPs to reach a slightly lower accuracy of around 87.5%. Similarly, on AIME, $g=4$ attains comparable peak accuracy to $g=1$, but with noticeably lower FLOPs cost across several configurations. This efficiency gain arises from fewer verifier invocations and reduced branching operations overall, as detailed in our cost model (Section 2.3). Moreover, the reduction in FLOPs effectively translates into decreased latency, as shown in Figure 1(c).
Rethinking the Optimal "Thinking Step". The observation that increasing $g$ (sparser verification) can improve accuracy while reducing compute budget is significant. This challenges the conventional verification granularity ($g=1$), where verification typically occurs at arbitrary delimiter boundaries (e.g., newlines), which may not always align with meaningful reasoning steps. First, as shown in Figure 1(a), consecutive PRM score differences are often negligible, indicating that the current definition of a “thinking step” is overly fine-grained. Second, a larger $g$ can be interpreted as defining a more substantial and semantically coherent "thinking step". Delaying verification and branching until the end of these extended segments may avoid injecting noise by evaluating incomplete reasoning fragments, thus enabling more effective search. Thus, in response to RQ1, we conclude that the current convention for verification granularity is sub-optimal, motivating a more flexible definition of verification granularity—beyond fixed token-level delimiters—and highlighting the need for a more sophisticated approach to optimizing $g$ (Section 5).
4.2 Ablation Studies on Verifier Quality and Model Scale
We conducted two ablation studies to investigate the impact of verifier quality and model scale on the optimal granularity, with results shown in Figure 4. To simulate a worst-case scenario for verifier quality, we replaced the standard verifier with a maximally noisy PRM that assigned a random score to each generation step. As hypothesized, when the verifier provides no useful signal, the optimal strategy consistently shifts to sparser verification ($g>1$), as frequent but unreliable guidance degrades search performance (Figure 4).
We validated our findings at different model scales by pairing various generators from the Qwen-2.5 series with Skywork-PRM verifiers of different sizes. The results (Figure 4) confirm that the optimal verification granularity remains dynamic across all tested scales.
Figure 4: Ablation on verifier quality and model sizes.
4.3 Trade-offs: Granularity, Verifier Parameter, and Branching
This section further explores how $g$ interacts with other key hyper-parameters: the verifier’s model parameter and the branch-out factor $B_{2}$ to optimize compute allocation under fixed total compute budgets. All experiments in this section utilize Qwen2.5-Math-7B as the generator.
Granularity vs. Verifier Parameter. We investigate whether it is more beneficial to use a strong verifier sparsely (larger $g$, larger verifier model) or a weaker verifier frequently (smaller $g$, smaller verifier model). Figure 5(a) compares two configurations: (1) $g=1$ (frequent verification) with a Small Verifier (Skywork-o1-1.5B), and (2) $g=2$ (sparser verification) with a Large Verifier (Skywork-o1-7B). As shown in Figure 5(a), at lower compute budgets, sparse verification with a larger verifier yields low accuracy. However, as the compute budget increases, employing a larger verifier more sparsely achieves higher peak accuracy. This suggests a clear preference for leveraging stronger, more discerning verifiers less frequently when sufficient computational resources are available.
(a) $g=1$, Small $\mathcal{V}$ VS $g=2$, Large $\mathcal{V}$
(b) Matched FLOPs with diff. $g$ and $B_{2}$
(c) Largest effective $g$ values
(d) Acc. diff. Majority Voting
(e) Acc. diff. Weighted Voting
Figure 5: (a) shows $g=2$ with a strong verifier outperforms $g=1$ with a weak verifier at high compute budget. (b) demonstrates limited performance gain from increasing branch factor $B_{2}$. (c) plots the largest effective $g$ across difficulty levels for different generators. (d),(e) show the accuracy gain of sparser verification with a stronger verifier across difficulty levels and number of generations.
Granularity vs. Branch-Out Factor Next, we explore the trade-off between the verification granularity $g$ and the branch-out factor $B_{2}$, at matched total compute FLOPs. Increasing $g$ means fewer verification calls and longer generation segments, while increasing $B_{2}$ means exploring more alternatives at each verification point. The verifier used in this experiment is Skywork-o1-1.5B. Figure 5(b) shows that configurations with a small granularity ($g=1,B_{2}=2$) with low branching tend to outperform other configurations at low or mid-range compute budgets. However, all configurations converge at very high compute. Comparing the above findings, a clear pattern emerges: investing the saved compute in the verifier parameter appears to be a more effective scaling strategy than simply increasing the number of branches.
4.4 Optimal Verification Granularity vs. Task Difficulty and Number of Samples
To assess the practical limits of sparse verification, we define the largest effective $g$ as the maximum verification granularity that retains at least 95% of the accuracy achieved at g=1g{=}1. We also analyze the accuracy gain of g=2g{=}2 over g=1g{=}1 across varying sample budgets and task difficulty levels. Figure 5(d),  5(e) show that the optimal $g$ is strongly influenced by both difficulty and sampling budget.
Increased Task Difficulty Demands Denser Verification. Across all models and compute budgets, we observe a consistent trend: as task difficulty increases, the largest effective $g$ decreases. For the hardest problems (Levels 4–5), g=1g{=}1 or g=2g{=}2 is often required to maintain performance at lower compute budget. Likewise, the advantage of g=2g{=}2 over g=1g{=}1 diminishes with difficulty, likely because harder tasks involve more reasoning errors, necessitating more frequent verifier intervention to stay on track.
Higher Number of Samples Tolerates Sparser Verification. In general, larger sample budgets allow for sparser verification, especially on easy to moderately difficult tasks. Higher budgets also amplify the performance gain of g=2g{=}2 over g=1g{=}1, as the generator can explore a wider search space, increasing the likelihood of finding a correct path despite less frequent verification.
5 Towards Adaptive Redefinition of the Verification Granularity
5.1 Adaptive Granularity Strategies
Our results (Figures 3, 5) show that the optimal $g$ in VG-Search depends on task difficulty, compute budget, and model capability. This motivates a more systematic approach to defining the search step, addressing RQ2. We formulate an optimization dual for selecting verification granularity: minimizing compute for a target accuracy vs. maximizing accuracy under a compute budget. For each, we propose a corresponding adaptive granularity strategy:
Strategy 1: Compute Minimization with Performance Parity (CM-$g$) CM-$g$ finds the largest $g^{*}$ that maintains accuracy within a tolerance $\epsilon$ of $g=1$, reducing compute. Given generator, difficulty $d$, and number of generations $n$: 1. Compute baseline accuracy $\text{Acc}(g{=}1,d,n)$. 2. Increase $g$ while $\text{Acc}(g,d,n)\geq\text{Acc}(g{=}1,d,n)-\epsilon$; return the largest $g^{*}$ satisfying the accuracy constraint.
Strategy 2: Accuracy Maximization with Budget Constraint (AM-$g$) AM-$g$ selects $g^{*}$ that maximizes accuracy under a fixed compute budget (as in number of generations): $g^{*}=\arg\max_{g\in\{1,\ldots,g_{\max}\}}\text{Acc}(g,d,n)$.
$n$ Metric Adaptive VG-Search Beam DVTS Best-of-N
Oracle Val Search
CM-$g$ AM-$g$ CM-$g$ AM-$g$
64 Acc. 89.2 90.1 88.6 90.1 87.0 89.2 89.5
FLOPs 5393 5844 5543 6145 12010 12010 11952
128 Acc. 90.5 90.6 89.9 90.1 89.3 89.7 89.3
FLOPs 15899 16200 11086 12290 24021 24021 23904
256 Acc. 90.7 90.8 90.4 90.4 88.8 89.5 89.2
FLOPs 21572 22775 28189 28189 48042 48042 47808
Table 1: Performance of adaptive granularity strategies (CM-$g$, AM-$g$) vs. baselines on MATH-500, using Qwen2.5-Math-7B as the proposer and Skywork-o1-1.5B as the verifier. Best accuracy and Lowest FLOPs are bold underlined. $n$ is the number of generations per question.
We compare against Beam Search, DVTS, and Best-of-N baselines on the MATH-500 test set. "val" indicates $g$ was tuned on a validation set MATH-250, while strategies under "oracle" use oracle $g$ selected on the test set. We set $\epsilon=0$ for CM-$g$. Table 1 shows that both AM-$g$ and CM-$g$ improve performance and efficiency. AM-$g$ consistently achieves higher accuracy, with gains up to 3.1% over Beam Search, while CM-$g$ provides substantial FLOPs savings, reducing compute by over 50% at $n=64$ and $n=256$ while maintaining or improving accuracy. For instance, CM-$g$ (val) at $n=128$ achieves 89.9% accuracy using just 11086 FLOPs—only 46% of the baseline budget. Although oracle-tuned variants perform slightly better, validation-tuned versions (CM-$g$ (val), AM-$g$ (val)) still outperform fixed-$g$ methods, demonstrating strong generalization and practicality. Under our cost model, $C_{\mathcal{G}}$ dominates, so most FLOPs savings come from pruning candidate paths during the Extend Step (Section 2.2). In summary, adapting verification granularity $g$ to task difficulty and compute budget offers a simple yet effective way, enabling more efficient and performant LLM reasoning.
5.2 Practicality and Convergence of Optimal Granularity Search
A key practical consideration is the overhead of finding the optimal granularity $g^{*}$. We analyzed its sample efficiency by selecting $g^{*}$ on validation subsets of increasing size and evaluating the performance on the MATH-500 test set. The results, shown in Table 2, demonstrate that the search converges rapidly. For both high ($n=256$) and medium ($n=64$) compute budgets, using just 5 validation samples achieves most of the possible accuracy gains over the baseline, confirming the minimal overhead of our adaptive approach.
This search constitutes a negligible, one-time cost that is amortized over all subsequent uses, ensuring our FLOPs comparisons are fair. The practical guideline is to reuse a pre-computed $g^{*}$ for a given model, task, and budget, and recompute it only when a core component, like the generator model, changes. For tasks with distinct subsets (e.g., varying difficulty levels), different $g^{*}$ values can be pre-computed and selected at inference time, making this a practical and effective alternative to the fixed-granularity convention.
Budget Strategy 0 samples 5 samples 100 samples 200 samples
$n=64$ CM-g 87.03% 88.57% 88.57% 88.57%
AM-g 87.03% 88.57% 89.19% 90.09%
$n=256$ CM-g 88.76% 90.37% 90.37% 90.43%
AM-g 88.76% 90.37% 90.37% 90.43%
Table 2: Test accuracy on MATH-500 vs. number of validation samples used to select $g^{*}$. Performance converges rapidly for both medium ($n=64$) and high ($n=256$) compute budgets.
6 Related Work
Test-Time Scaling (TTS). Recent work on scaling test-time compute can be broadly categorized into two approaches [43, 19]: i) Internal scaling and ii) sampling-based TTS. Internal scaling methods enable models to scale their test-time compute by adapting their parameters through Supervised Fine-tuning (SFT) [42, 24] or Reinforcement Learning (RL) [33, 23, 41]. However, without explicit compute budget control [1, 3], these approaches often result in unnecessarily long CoT, even for simple queries, leading to inefficient inference costs. Sampling-based TTS, by contrast, dynamically allocates test-time computation during deployment, without requiring additional tuning to modify the model’s parameters offline. Sampling-based TTS [10, 32, 29, 9], orthogonal to internal scaling approaches, allows for more fine-grained control over compute usage and can further improve algorithmic performance. As noted by [32, 22], multi-step lookahead approaches, such as Monte Carlo Tree Search (MCTS) [14], adaptive thought rollback [8], and hierarchical multi-step verification [36], incur substantial sampling overhead during inference, we do not consider such methods in this work.
Verification-Guided Search. Verifying the quality of generated outputs to guide sampling-based TTS has been shown to significantly improve the efficiency of test-time computation [22, 32, 4]. Based on the granularity of verification, existing approaches can be broadly classified into two categories: (i) outcome verification and (ii) process verification. Outcome verification [10] evaluates the entire response after the generation of a complete sample. This approach is widely used in sampling-based TTS, such as Best-of-N strategy [9], where a verifier selects the most promising answer from a set of full samples. While simple and effective, outcome verification provides no intermediate feedback, which can lead to wasted computation on low-quality completions. Process verification [21, 44], often implemented as a process reward model (PRM), evaluates intermediate reasoning steps during generation. These models assign scores to partial outputs, allowing the inference process to be guided. PRMs have been shown to improve sample efficiency and enable early pruning of unpromising token paths, leading to more compute-efficient search. Despite their promise, both outcome and process verification methods treat the verifier as a fixed component and typically apply it at a pre-defined granularity. In contrast, our work introduces an adaptive approach that dynamically optimizes verification granularity to maximize performance. Due to the verification inefficiency of Generative PRMs [44] under matched compute budgets [31], we leave their investigation to future work.
Compute Optimal Scaling. Previous research [32] indicates that the optimal choice between Best-of-N and Beam Search often depends on factors such as the available compute budget and the problem characteristics. Yet, restricting selection to these two discrete options limits flexibility and expressiveness in identifying an optimal compute scaling strategy. By introducing variable granularity ($g$), VG-Search spans a continuous spectrum of search behaviors between these extremes, offering potentially greater expressiveness for the optimal TTS strategy.
Prior work has investigated how test-time strategies, compute budgets, and problem difficulty interact, showing that optimal allocation of inference-time compute can be more effective than merely scaling model parameters [32]. Scaling the number of verification attempts per candidate has been shown to improve verification accuracy and overall task performance [30]. Recent work [31] also explores compute-optimal strategies for deciding when to generate new solutions versus verifying existing ones, taking into account the number of verification and output samples. [20] demonstrates that employing multiple diverse verifiers can enhance performance, as different verifiers are likely to detect different types of errors. [7] proposes a consistency-based model switching method that leverages multiple generator models to efficiently scale test-time compute, showing that a diverse generator pool can more effectively explore the solution space. [8] introduces adaptive thought rollback, but does not consider verification granularity. Despite the progress, optimally adapting the verifier’s invocation frequency during test-time scaling remains a key unanswered question.
7 Conclusion
This paper investigates the critical role of the verification granularity $g$ in verifier-guided test-time scaling, introducing VG-Search to facilitate exploration. Our experiments show that the optimal verification granularity is highly dynamic—depending on generator capability, task difficulty, and compute budget. Building on these insights, we challenge the current convention and propose adaptive strategies for selecting $g$ to optimize efficiency and performance. This work motivates future research that explores more advanced approaches in optimizing the verification granularity $g$.
Appendix A Appendix
A.1 Optimal Compute Allocation between Proposer and Verifier
Optimizing the allocation of computational resources between proposers and verifiers is crucial for maximizing overall system performance. While previous work [31] under a fixed compute budget has primarily focused on varying the ratio of samples generated by the proposer to those evaluated by the verifier, the impact of asymmetric model capabilities—specifically, how distributing model parameters differently between the proposer and verifier affects performance—has remained largely unexplored. This section addresses this gap by investigating the optimal allocation of model parameters between these components, while adhering to a constant total compute budget, to determine how performance scales with their relative complexities.
We evaluate two configurations: Small Proposer, Large Verifier (S-P, L-V) and Large Proposer, Small Verifier (L-P, S-V), using DVTS and Beam Search. Performance is measured by accuracy against the number of samples ($n$) on tasks of varying difficulty (Level 1, Level 5) and averaged (Figure 6).
(a) DVTS
(b) Beam Search
Figure 6: Comparison of Proposer-Verifier configurations under (a) DVTS and (b) Beam Search strategies. Accuracy (%) is plotted against the number of samples ($n$) for Level 1, Level 5, and Average Accuracy. Blue line: Small Proposer, Large Verifier (S-P, L-V). Orange line: Large Proposer, Small Verifier (L-P, S-V).
Key observations include:
Verifier Quality Dictates Scaling: With both DVTS and beam search (Figure 6), L-P, S-V starts strong on easier tasks but scales poorly with more samples ($n$), especially on harder tasks, suggesting that more samples without effective verification is counterproductive.. S-P, L-V shows more robust scaling, indicating a strong verifier better utilizes increased compute for broader exploration. A weak verifier bottlenecks the scaling.
Strategy Determines Optimal Configuration: The best Proposer-Verifier pairing is strategy-dependent. For DVTS, S-P, L-V generally achieves higher peak performance or scales better for complex tasks. For Beam Search (Figure 6(b)), L-P, S-V consistently and significantly outperforms. DVTS’s diverse exploration of the solution space benefits more from a strong verifier, making a simpler proposer adequate.
Task Difficulty Modulates Preference: On easier tasks (Level 1), S-P, L-V can be competitive, probably because the proposer’s performance is not a bottleneck for easier tasks. However, for harder tasks (Level 5), a strong proposer becomes more critical for beam search.
These results highlight that optimal compute allocation between proposer and verifier is not fixed but depends on the search strategy, task complexity, and available budget. The observation motivates more advanced compute allocation between verifiers and proposers.
A.2 FLOPs Calculation using the Cost Model
Following prior work [31], for discriminative PRMs, we assume that $C_{\mathcal{V}}=2P_{v}$, since the verifier outputs a single score per evaluation. Under this assumption, Equation 3 simplifies to:
$C_{\text{total}}=2\cdot B_{1}\cdot\frac{L}{g}\cdot\left[T\cdot(g-1+B_{2})\cdot P_{g}+B_{2}\cdot P_{v}\right],$ (4)
where $L$ is the average solution length in generation steps, $T$ is the average number of tokens per generation step. To simplify compute tracking across experiments, we define a normalized compute proxy:
$C(g,B_{1},B_{2})=\frac{\lambda\cdot B_{1}\cdot(g-1+B_{2})+B_{1}\cdot B_{2}}{g},$ (5)
where $\lambda=\frac{T\cdot P_{g}}{P_{v}}$ captures the relative cost of generation versus verification. This proxy is proportional to the actual total compute budget and is used throughout all experiments in the paper.
A.3 Experiment Details for VG-Search
Experiment Setup. To investigate the optimal $g$ for a given difficulty level and computational budget, we conduct experiments assessing both the oracle and the validation-tuned accuracies for the CM-$g$ and AM-$g$ strategies. We use Qwen2.5-Math-7B [40] as proposer, and Skywork-o1-1.5B [27] as verifier. The number of generations is set to $n\in\{4,16,64,128,256\}$, with a Branch-Out Factor $B_{2}=4$, temperature $0.8$, and Top-p $1.0$. For each generation step, the maximum token number is set to $2,048$, employing Last scoring and majority voting. The number of iterations $I$ corresponds to $g$: for $g\in\{1,2,3,4\}$, we set $I\in\{12,6,4,3\}$ to ensure equal compute budget. The prompt template from Qwen [40] is adopted. DVTS, Best-of-N (BoN), and Beam Search are used as baselines, configured with the same parameters as VG-Search.
[…]
For ideal accuracy, we search for optimal $g$ in test dataset for every combination of number of generations and difficulty level, and compute two accuracies for each number of generation: unweighted and weighted, the former corresponds to the dataset having equal number of questions in different difficulty levels, and the latter corresponds to actual difficulty distribution in the dataset. For the searched accuracy, the only difference is that we search $g$ in the validation dataset, and obtain the accuracy in the test dataset for searched $g$.
Baselines and Voting Methods. Our codebase is based on HuggingFace’s search-and-learn [4]. We adopt three mainstream TTS strategies: Best-of-N [5], Beam Search [32], and DVTS [4]. For all strategies we explore the number of samples of 4, 16, 64, 128, and 256. For Beam Search and DVTS, we use a fixed Branch-Out Factor $B_{2}=4$ following [32]. For all strategies, we use majority voting [37]. Temperature is set to 0.8 for all problems following HuggingFace’s setup.
Hardware and Frameworks. We conduct our experiments on NVIDIA H100 [12] and A100 [11] GPUs, using CUDA 12.8 on Ubuntu 22.04. The vLLM library (v0.6.3) [18] is employed for model execution. The experiments for VG-Search on MATH-500 take from 1 to 5 hours, depending on the candidate number. To obtain accurate system-level runtime measurements, we use Nsight Systems [26] and employ the NVIDIA NVTX [13] extension to categorize execution time across different models. For the latency measurement in Figure 1(c), we use Qwen2.5-Math-1.5B as generator, while using RLHFlow/Llama3.1-8B-PRM-Deepseek-Data as the verifier.
A.4 Evaluation with an Ensemble Verifier
To broaden the scope of our evaluation, we conducted additional experiments using a multi-verifier ensemble. This approach can enhance verification quality by aggregating signals from diverse models [20]. For these experiments, we used Qwen-2.5-Math-1.5B as the generator. The verifier was an ensemble of two different Process Reward Models (PRMs): Skywork-1.5B [27] and RLHFlow’s Llama3.1-8B-PRM [39]. The final verification score for each reasoning step was computed by averaging the scores from these two models.
The results, presented in Table 3, align with and reinforce the main conclusion of our paper. We observe that the optimal verification granularity ($g$) is dynamic and highly dependent on the compute budget (i.e., the number of generations, $n$). While dense verification ($g=1$) is sometimes preferable at low sample counts, sparser verification ($g>1$) consistently achieves the highest accuracy on both the MATH-500 and AIME datasets as the compute budget increases. This confirms that our findings on adaptive verification granularity generalize even when using a stronger, ensembled verifier.
n MATH-500 (%) AIME (%)
g=1 g=2 g=3 g=4 g=1 g=2 g=3 g=4
4 76.0 73.2 71.6 69.6 12.22 8.89 11.11 13.33
16 80.0 77.4 76.4 78.4 13.33 13.33 16.67 14.44
64 79.4 80.4 80.8 81.0 16.67 16.67 16.67 18.89
128 80.8 82.2 82.2 80.8 15.56 24.44 23.33 18.89
256 80.2 80.0 80.2 81.2 23.33 23.33 23.33 23.33
Table 3: Accuracy of VG-Search with an ensemble verifier across MATH-500 and AIME datasets. For each number of generations ($n$), the highest accuracy in each row is bolded, indicating the performance achieved at the optimal granularity $g$.
A.5 Applicability to Tasks with Outcome-Level Feedback
VG-Search is applicable to tasks with outcome-level feedback, provided a PRM can be trained on intermediate generation steps. In such scenarios, the core search algorithm remains unchanged. However, for open-ended generation tasks like code synthesis where outputs are diverse, standard majority voting is not feasible for final answer selection. A practical alternative is to select the single candidate trajectory with the highest cumulative or final PRM score.
To demonstrate this, we evaluated VG-Search on the HumanEval benchmark using LLaMA-3.2-1B as the generator and Skywork-o1-1.5B-PRM as the verifier. We used newline characters to delimit generation steps. The results, reported as pass@1 scores in Table 4, show that sparser verification ($g>1$) can outperform dense verification ($g=1$), particularly as the compute budget ($n$) increases. This confirms that our central conclusion—that the optimal verification granularity is dynamic—generalizes to open-ended generation tasks.
n g=1 g=2 g=3 g=4
4 0.250 0.274 0.238 0.220
16 0.335 0.305 0.268 0.311
64 0.256 0.335 0.274 0.305
Table 4: Pass@1 scores on HumanEval using VG-Search, selecting the highest-scoring candidate. The best performance for each compute budget ($n$) is bolded.
A.6 Standard Deviation with Multiple Trials
To evaluate the robustness of our results, we conducted experiments over multiple trials using different random seeds. This analysis provides a more rigorous assessment of variance and confirms the stability of our conclusions. The following results were generated using the Qwen-Math-7B generator and the Skywork-o1-1.5B PRM on the MATH-500 dataset, averaged over three distinct runs.
Table 5 shows the mean accuracy and standard deviation for this experiment. The results from multiple trials are consistent with our original findings, demonstrating the stability of the observed trends. The low standard deviations across all configurations reinforce this conclusion. While dense verification ($g=1$) performs well at smaller compute budgets, sparser verification ($g>1$) consistently achieves higher accuracy as the number of samples increases ($n\geq 64$).
n g=1 g=2 g=3 g=4
4 84.13 $\pm$ 0.50 82.27 $\pm$ 0.24 81.73 $\pm$ 0.39 80.40 $\pm$ 0.43
16 86.80 $\pm$ 0.28 86.13 $\pm$ 0.22 86.13 $\pm$ 0.25 86.40 $\pm$ 0.16
64 86.20 $\pm$ 0.33 87.27 $\pm$ 0.15 88.27 $\pm$ 0.34 87.87 $\pm$ 0.19
128 86.67 $\pm$ 0.19 87.60 $\pm$ 0.57 88.27 $\pm$ 0.21 88.47 $\pm$ 0.32
256 86.67 $\pm$ 0.19 87.93 $\pm$ 0.34 89.13 $\pm$ 0.27 88.73 $\pm$ 0.19
Table 5: Mean accuracy $\pm$ standard deviation over 3 runs with different seeds on MATH-500. The highest mean accuracy for each sample count ($n$) is bolded.
A.7 Limitations
While our proposed adaptive strategies to optimize the verification granularity $g$ based on overall problem characteristics like difficulty and compute budget, a key limitation is that the chosen $g$ remains fixed throughout the generation process for a single problem instance. This approach does not account for potential variations in reasoning complexity or generator confidence within different stages of solving a single problem. A more sophisticated approach, potentially offering further efficiency gains, would involve dynamically adjusting the granularity $g$ on a step-by-step basis during inference, adapting verification frequency based on real-time signals. Developing such fine-grained, intra-problem adaptive granularity remains an important direction for future research.
A.8 Case Study of VG-Search
Figure 7: A case study of VG-Search. $B_{1}=4,B_{2}=2,g=3$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction accurately reflect the paper’s contributions, evidenced by the following sections.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Appendix A.7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have provided our code as the supplementary material, as well as experiment details and instructions.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided our code as the supplementary material, as well as experiment details and instructions.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided our code as the supplementary material, as well as experiment details and instructions. We also describe experiment details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Our main results do not contain randomness as we fixed the seed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix A.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have made the citations.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: They are well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not use LLMs for method development.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.